OpenStudy (laddiusmaximus):

Find a point c satisfying the conclusion of the Mean Value Theorem for the following function and interval. f(x)=x^(-2) [1,5]

OpenStudy (laddiusmaximus):

I know to use f(b)-f(a)/b-a becomes (1/5-1/1)/(4)

OpenStudy (laddiusmaximus):

which equals -16/5

OpenStudy (laddiusmaximus):

im lost after this

OpenStudy (zarkon):

\[f(5)=5^{-2}=\frac{1}{25}\]

OpenStudy (zarkon):

you need to solve \[\frac{f(5)-f(1)}{5-1}=f'(c)\] for \(c\)

OpenStudy (laddiusmaximus):

still lost

OpenStudy (experimentx):

find dy/dx ... put it to the value calculated from above and solve for x

OpenStudy (experimentx):

dy/dx = ((1/25) - 1)/(5-1) ... solve for x

OpenStudy (experimentx):

c = http://www.wolframalpha.com/input/?i=-2x%5E-3+%3D+-24%2F100

OpenStudy (experimentx):

2.0374 (( [1, 5]

OpenStudy (anonymous):

\[ \frac{f[5]-1}{5-1}= f'[c] \] \[ -\frac{6}{25}= -\frac 2 {c^3} \] \[3 c^3= 25 \] \[c = (\frac {25}3)^{1/3}\]